OpenStudy (nbuzhduga):

Solve the following system of equations: (1 point) −2x + y = 1 −4x + y = −1

OpenStudy (yoloshroom):

This is a homework review problem, right? <3

OpenStudy (anonymous):

x=1,y=3

OpenStudy (yoloshroom):

You need to explain.

OpenStudy (anonymous):

oh..... -2x+y=1 -4x+y=-1 y=1+2x -4x+1+2x=-1 -2x=-1 x=1

OpenStudy (anonymous):

You need to isolate the x and isolate the y to one side so you can solve. You can always set the problem equal to zero and plug numbers in to check them

OpenStudy (anonymous):

and next we put x on one of equations -2+y=1 y=3

OpenStudy (retireed):

It is easier to subtract equation 2 from equation 1. −2x + y = 1 −4x + y = −1 4x - y = 1 sometimes it is easier to multiple equation 2 by - 1 and add in either case the y is eliminated 2x + 0 = 2 solve for x and plug it back in and solve for y